Hello! I need some assistance on how to find a missing side of a rectangle. The question is: Mary drew a model of her school, the length of the school is 27ft and the width is 84ft. The model's width is 3ft and the length is "x". What length in feet, is x? I'm not completely sure on how to begin this problem, or at least how to set up an equation to try to solve this problem.If you could please help on how to solve this problem, it would be most appreciative. Thank You!
Answer by stanbon(75887)

Mary drew a model of her school,
the length of the school is 27ft
and the width is 84ft.
--------------------------------------------
The model's width is 3ft and the length is "x".
------------------------------------------------
What length in feet, is x?
model length/length = model width/width
-------------------------------------------
x/27 = 3/84
--------
x = 27(3/84)
---
x = 81/84 = 0.96 ft
=================================
Cheers,
Stan H.